OpenStudy (anonymous):

what is the derivative of 6^(lnx^2)

OpenStudy (anonymous):

\[\ln x^2 = 2lnx\]

OpenStudy (anonymous):

so, yours y = 36^lnx. take ln both sides lny = ln 36^lnx = lnx * ln 36

OpenStudy (anonymous):

take derivative is much more easier than original one. hope that helps

OpenStudy (anonymous):

If you meant to say \[y=6^{(\ln x)^2},\] \[\ln y=(\ln x)^2\ln6\\ \frac{1}{y}y'=2\ln6(\ln x)\frac{1}{x}\\ y'=\ln36\frac{\ln x}{x}\cdot y\\ y'=\ln36\frac{\ln x}{x}\cdot 6^{(\ln x)^2}\]